OpenStudy (anonymous):

Which value is a solution for the equation cot(x/2)=1? pi/4 3pi/4 7pi/4 5pi/2

OpenStudy (agreene):

cot(x/2)=1 cot^(-1) both sides x/2=pi/4 x= pi/2 since it is cot and thus osculating, we can say: \[x=\frac{n\pi}{2} : n\in\mathbb{Z}\] so 5pi/2 is a solution.